The goal of improving information seeking and retrieval is currently at its zenith both in industry and academia. However, there are still many challenges to be met towards the goal of innovating to improve search for everyone, in every applicable context. This talk will review a set of developing ideas and some ongoing research including personalization, filtering, user-defined metadata and information architecture that provide both research and development opportunities with potential impact to advance search utility.